# Advancing Semantic Textual Similarity Modeling: A Regression Framework with Translated ReLU and Smooth K2 Loss

### v1

@@ 4.1 STS Performance without Task-specific Training Data
4.2 STS Performance with Task-specific Training Data
Although the Siamese neural network, augmented by our regression framework and innovative loss functions, has exhibited significant performance enhancements, a disparity persists relative to prevailing contrastive learning methods. To bridge this gap, we exploit another critical advantage of the Siamese architecture: its capacity to fully utilize task-specific training data.
Among the seven STS benchmarks (STS12-16, STS-B, and SICK-R), STS-B and SICK-R come with their own training datasets. The STS-B training set comprises 5,749 sentence pairs with similarity scores ranging from 0 to 5, whereas the SICK-R training set includes 4,500 pairs, scored from 1 to 5. To standardize these scores, we apply a transformation to each sample in the SICK-R dataset. Subsequently, we concatenate these two sets and round all sample labels to integers, resulting in a task-specific training dataset containing 10,249 sentences pairs. While the sample quantity provided by this newly introduced dataset is approximately only one percent of the NLI corpus, combining them has been sufficient for us to surpass leading contrastive learning approaches.
Continuing from the checkpoint established in Section 4.1, we further fine-tune our model using this compact, task-specific dataset with Smooth K2 Loss. Adhering to our protocol, checkpoints are preserved based on STS-B development set performance. The updated results across the seven STS benchmarks are summarized in Table 3, illustrating an improvement in our method’s average Spearman correlation for BERT and RoBERTa from 76.03 and 76.04 to 82.93 and 83.23, respectively. These outcomes exceed those achieved by leading contrastive learning methods, such as SimCSE, PromptBERT, Jina Embeddings 2 (Günther et al., 2023), and Nomic Embed (Nussbaum et al., 2024), and set new SOTA performance.
@@ 4.2 STS Performance with Task-specific Training Data
| Contrastive Pre-training Model | ||||||||
| Jina Embeddings v2 base | 74.28 | 84.18 | 78.81 | 87.55 | 85.35 | 84.85 | 78.98 | 82.00 |
| Nomic Embed Text v1 | 65.19 | 81.67 | 74.00 | 83.58 | 81.87 | 76.43 | 75.41 | 76.88 |
| Implementation on BERT | ||||||||
| SimCSE BERT | 75.30 | 84.67 | 80.19 | 85.40 | 80.82 | 84.25 | 80.39 | 81.57 |
@@
| Ours STS-B SICK-R train | 73.68 | 88.42 | 86.10 | 86.56 | 79.63 | 84.12 | 82.01 | 82.93 |
| Implementation on RoBERTa | ||||||||
| SimCSE RoBERTa | 76.53 | 85.21 | 80.95 | 86.03 | 82.57 | 85.83 | 80.50 | 82.52 |
| PromptRoBERTa | 76.75 | 85.93 | 82.28 | 86.69 | 82.80 | 86.14 | 80.04 | 82.95 |
| Ours STS-B SICK-R train | 73.83 | 89.00 | 84.16 | 87.95 | 81.94 | 84.64 | 81.07 | 83.23 |
@@
Contrastive learning methods, by contrast, are generally unable to leverage the detailed, multi-level annotated information provided by STS datasets. The prevalent contrastive learning loss function, InfoNCE Loss, serves as an illustrative case for this limitation. For any input sentence , InfoNCE Loss computes the similarity between its encoding and that of its positive instance in the numerator, while the denominator aggregates similarity calculations between and encodings of other samples within the same batch, aiming to draw similar samples closer and push dissimilar ones apart. The standard formulation of InfoNCE Loss, where represents the batch size and a temperature hyperparameter, is as follows:
While this mechanism effectively refines the semantic space distribution of PLMs, it is constrained to utilizing only text pairs with the highest similarity ratings. Since InfoNCE Loss merely includes numerator and denominator components, it distinguishes only whether two texts are similar or not. Given the denominator is composed of other samples within the same batch, the only part that can be filled with labeled data is the numerator.
In contexts where more detailed, domain-specific data is available, the shortcomings of contrastive learning in not being able to effectively harness multi-level label information, only performing coarse semantic distinctions, becomes more evident. A potential pathway is to combine our regression framework with contrastive learning. By supplementing a contrastively trained model with our Siamese neural network architecture, it may be possible to capture finer semantic nuances. This avenue of exploration holds promise for future work, potentially enhancing the applicability and efficacy of our approach.
4.3 Performance under Different Hyperparameter Settings
@@
In this study, we introduce two innovative loss functions, Translated ReLU and Smooth K2 Loss, each characterized by two critical hyperparameters: and . The parameter primarily controls the gradient of the loss function, while sets the tolerance threshold for model predictions. To discern the influence of these hyperparameters on model performance, we conduct a series of experiments during both the initial training phase with NLI datasets and the subsequent fine-tuning phase with task-specific training data. The outcomes of these investigations are consolidated in Tables 4. Rather than executing an exhaustive grid search, initial values are selected based on our preliminary insights, followed by incremental adjustments. This implies that there may still be room for further improvement in our model’s performance.
@@
| Without task-specific training data | ||||
@@
| With task-specific training data | ||||
| BERT | Smooth K2 Loss | 4 | 0.2 | 82.89 |
| BERT | Smooth K2 Loss | 3.5 | 0.25 | 82.89 |
| BERT | Smooth K2 Loss | 4 | 0.3 | 82.90 |
| BERT | Smooth K2 Loss | 4 | 0.25 | 82.93 |
| RoBERTa | Smooth K2 Loss | 4 | 0.3 | 82.86 |
| RoBERTa | Smooth K2 Loss | 3.5 | 0.25 | 82.90 |
| RoBERTa | Smooth K2 Loss | 3.5 | 0.3 | 83.18 |
| RoBERTa | Smooth K2 Loss | 3 | 0.25 | 83.23 |
The experimental results from Table 4 reveal minor fluctuations in model performance across diverse hyperparameter configurations, which affirms the resilience and robustness of our proposed methodology. This stability highlights the inherent adaptability of our regression framework as well as loss functions, suggesting their applicability across a wide range of modeling scenarios without necessitating extensive hyperparameter optimization.
@@ 4.4 Ablation Studies
In Section 4.1, we initially demonstrate the effectiveness of our regression framework by comparing the performance differences of models utilizing both classification-based and regression-based strategies for STS tasks. Then, we elucidate the significance of zero-gradient buffer zones by comparing the performance of models when selecting Translated ReLU or L1 Loss, and Smooth K2 Loss or MSE Loss as the loss function. These comparisons directly align with the three core innovations of this paper and fulfill the role of ablation experiments.
Here, we enhance our ablation study with an evaluation of our network architecture as depicted in Figure 1. Specifically, we aim to determine the necessity of concatenating , , and their element-wise difference in the final linear layer of the model. For this purpose, we employ both BERT and RoBERTa models under the same experimental conditions outlined in Section 4.1, with results detailed in Table 5. The findings indicate that the concatenation method is the most effective for both PLMs, thus further validating the rationality of our proposed scheme.
@@ 5 Conclusion
In this paper, we propose an innovative regression framework accompanied by two simple yet efficacious loss functions: Translated ReLU and Smooth K2 Loss, to address multi-category STS tasks. Compared to traditional classification strategies, our regression framework achieves superior performance while reducing the parameter count of the model’s output layer. Further empirical evidence demonstrates that when supplemented with task-specific training data, our approach can surpass prevailing contrastive learning methods, achieving state-of-the-art performance on seven STS benchmarks.
@@
Due to the lack of suitable baselines and limited computational resources, the experiments in this paper are primarily centered on the discriminative PLMs such as BERT and RoBERTa, rather than recently advanced generative models (e.g., LLaMA (Touvron et al., 2023)). However, it is important to note that, compared to generative PLMs, BERT possesses a much smaller parameter count, which leads to higher inference efficiency. This attribute is particularly valuable in large-scale information retrieval scenarios.

### v2

@@ 4.1 STS Performance Based on Traditional Discriminative Pre-Trained Models
Our experimental setup here closely mirrors that of Sentence-BERT, leveraging fine-tuning on BERT or RoBERTa with a composite corpus derived from the SNLI and MNLI datasets. These NLI datasets categorize sentence pairs into three distinct classes: contradiction, neutral, and entailment. Sentence-BERT maps these classes to 0, 2, and 1, respectively, and employs a classification strategy for training (Reimers and Gurevych, 2019). In contrast, our method sequentially maps contradiction, neutral, and entailment to 0, 1 and 2. This mapping reflects the natural order of semantic similarity, from least to most similar, thereby enabling our regression framework to better capture the progressive relationships between categories.
| PLM | Loss | ||
| BERT | Translated ReLU | 2.5 | 0.25 |
| BERT | Smooth K2 Loss | 2 | 0.25 |
| RoBERTa | Translated ReLU | 1 | 0.25 |
| RoBERTa | Smooth K2 Loss | 3 | 0.25 |
For computational efficiency, we uniformly set the batch size to 16 and limit training to a single epoch, with model checkpoints saved based on performance metrics on the STS-B development set. The specific hyperparameter settings for Translated ReLU and Smooth K2 Loss are detailed in Table 2. During evaluation, we assess the model’s average Spearman correlation across seven STS tasks via the SentEval toolkit (Conneau and Kiela, 2018). The results of these experiments are summarized in Table 1, from which we distill insights along three pivotal aspects:
1. Classification Strategy vs. Regression Strategy: Our regression framework, particularly when utilizing Smooth K2 Loss, yields an average Spearman correlation of 76.03 for BERT and 76.04 for RoBERTa. These figures significantly outstrip those attained through Sentence-BERT and the classification strategy with Cross-Entropy Loss, highlighting the regression-based modeling’s superiority in both reducing the output layer’s parameter size and enhancing semantic discrimination in multi-category STS tasks.
2. Efficacy of the Zero-Gradient Buffer Zone: The adoption of Translated ReLU improves performance for both BERT and RoBERTa beyond what is achieved with L1 Loss. Similarly, employing Smooth K2 Loss surpasses MSE Loss on both PLMs. These comparisons underline the benefit of incorporating a zero-gradient buffer zone, which helps balance the model’s attention across diverse samples in regression-modeled multi-category classification tasks.
@@
| CT-SBERT | 74.84 | 83.20 | 78.07 | 83.84 | 77.93 | 81.46 | 76.42 | 79.39 |
| SimCSE-BERT | 75.30 | 84.67 | 80.19 | 85.40 | 80.82 | 84.25 | 80.39 | 81.57 |
@@
| PromCSE-BERT | 75.58 | 84.33 | 79.67 | 85.79 | 81.24 | 84.25 | 80.79 | 81.81 |
| Nomic Embed Text v1 | 73.75 | 85.03 | 80.52 | 87.40 | 83.55 | 83.90 | 76.52 | 81.52 |
| Nomic Embed Text v1 + Contrast | 76.10 | 85.79 | 80.58 | 87.35 | 83.54 | 85.16 | 72.33 | 81.55 |
| Nomic Embed Text v1 + Ours | 73.06 | 86.63 | 81.06 | 87.67 | 83.43 | 85.18 | 82.75 | 82.83 |
| Jina Embeddings v2 | 74.28 | 84.18 | 78.81 | 87.55 | 85.35 | 84.85 | 78.98 | 82.00 |
| Jina Embeddings v2 + Contrast | 76.04 | 86.37 | 80.16 | 86.53 | 85.24 | 84.31 | 74.18 | 81.83 |
| Jina Embeddings v2 + Ours | 75.17 | 86.10 | 79.96 | 88.44 | 85.01 | 86.83 | 83.34 | 83.55 |
3. Adaptive Gradients Aligned with Prediction Errors: Models trained with Smooth K2 Loss outperform those utilizing Translated ReLU, and models employing MSE Loss exceed those with L1 Loss. This evidences the advantages of dispensing differentiated gradients in line with prediction-ground truth deviations, especially when leveraging high-quality datasets like NLI.
Collectively, these findings substantiate the merit of (1) adopting a regression framework for multi-class STS tasks and (2) enhancing traditional regression loss functions with a zero-gradient buffer zone to optimize model performance.
4.2 STS Performance Based on Contrastive Learning Pre-Trained Models
While the Siamese neural network, augmented by our regression framework and innovative loss functions, has exhibited significant performance improvements, a gap remains when compared to leading contrastive learning methods. To address this, we exploit the strengths of Siamese architectures in fully utilizing annotated data and explore whether it can be combined with top-performing contrastive learning models.
Jina Embeddings v2 (Günther et al., 2023) and Nomic Embed (Nussbaum et al., 2024) are two recently released embedding models that employ multi-stage contrastive learning strategies during pre-training, combining supervised and unsupervised approaches to optimize the networks. Both have achieved state-of-the-art results on the MTEB leaderboard (Muennighoff et al., 2023). Therefore, if our method can further enhance the performance of these models, it would provide valuable insights for future research.
Among the seven STS benchmarks (STS12-16, STS-B, and SICK-R), STS-B and SICK-R come with their own training datasets. Specifically, STS-B contains sentence pairs with similarity scores ranging from 0 to 5, while SICK-R includes pairs with scores from 1 to 5. To ensure accurate evaluation, we performed strict data filtering to remove any training text pairs that appeared in the test sets. Details of this filtering process are provided in Appendix A. We then applied a linear transformation, , to convert all SICK-R training labels to the range [0, 5] and merged them with the filtered STS-B training set. This procedure resulted in a fine-grained, task-specific corpus containing 5,398 sentence pairs.
Since Jina Embeddings v2 and Nomic Embed have undergone pre-training on massive texts, their model parameters have favorable initial distributions. In contrast, our newly introduced linear layer is randomly initialized (Figure 1). To facilitate effective joint training, we first freeze the entire PLM and only update the linear layer using the NLI dataset described in section 4.1. After completing this step, we optimize both the PLM and the linear layer with the filtered STS training data. A schematic diagram of this workflow is shown in Figure 3. Throughout the entire procedure, Smooth K2 Loss is employed as the loss function.
The performance of Nomic Embed and Jina Embeddings v2 on the seven STS tasks before and after fine-tuning is presented in Table 3. The results demonstrate that our network framework effectively enhances the performance of both models and surpasses BERT-based methods with comparable parameter sizes. Notably, we also test the impact of further updating the PLM using contrastive learning, which requires additional processing of the 5,398 training samples obtained earlier. To illustrate this, we take InfoNCE Loss (Oord et al., 2018), the most widely adopted contrastive learning loss function, as an example.
@@ 4.2 STS Performance Based on Contrastive Learning Pre-Trained Models
For any input sentence , InfoNCE Loss computes the similarity between its encoding and that of its positive instance in the numerator, while aggregating the similarity calculations between and other samples within the same batch in the denominator. This formulation aims to bring similar samples closer and push dissimilar ones apart. Equation 3 presents the standard expression of InfoNCE Loss, where represents the batch size and denotes a temperature hyperparameter.
As indicated by Equation 3, the only component of InfoNCE Loss that can be filled with labeled data is the similarity calculation between positive samples in the numerator. Consequently, contrastive learning is limited to utilizing only text pairs with the highest similarity ratings. To work within this constraint, we selected 1,543 samples from the 5,398 training pairs by adopting a threshold of 4.0 to filter out positive sample pairs. As it can be observed in Table 3, after discarding such a large portion of annotation information, contrastive learning yields little improvement and may even lead to model collapse, causing performance degradation. In contexts where more detailed, domain-specific data is available, the shortcomings of contrastive learning in not being able to effectively harness multi-level label information, only performing coarse semantic distinctions, becomes more pronounced.
4.3 Computational Resource Overhead
In addition to its inability to fully leverage fine-grained annotated data, the high memory requirements of contrastive learning also pose a challenge for many researchers. In this section, we compare the computational resource consumption of our method with that of SimCSE during training, based on four 24GB NVIDIA GPUs. The results are summarized in Table 4, where both BERT and RoBERTa are the base versions.
Despite setting the maximum sequence length for SimCSE to approximately 40% of our method’s default configuration, its memory usage remains significantly higher, reaching an astonishing 81GB. Thus, overall, our Siamese neural network strategy is more suitable for resource-constrained environments.
| PLMs | Method | Length | Memory |
| BERT | SimCSE | 100 | 81.30 GB |
| Ours | 256 | 41.27 GB | |
| RoBERTa | SimCSE | 100 | 81.61 GB |
| Ours | 256 | 42.33 GB |
4.4 Impact of Different Hyperparameter Settings
In this study, we introduce two novel loss functions, Translated ReLU and Smooth K2 Loss, each characterized by two critical hyperparameters: and . The parameter primarily controls the gradient of the loss function, while defines the tolerance threshold for model predictions. To discern the influence of these hyperparameters on model performance, we conducted a series of experiments across both traditional discriminative PLMs (BERT, RoBERTa) and the latest contrastive learning PLMs (Nomic Embed v1, Jina Embeddings v2).
@@ 4.4 Impact of Different Hyperparameter Settings
The outcomes of these investigations are consolidated in Tables 5. Rather than executing an exhaustive grid search, initial values were selected based on our preliminary insights, followed by incremental adjustments. This implies that there may still be room for further improvement in our model’s performance.
@@
| Implementation on Traditional Discriminative PLMs | ||||
@@
| Implementation on Contrastive Learning PLMs | ||||
| Nomic v1 | Smooth K2 Loss | 3.5 | 0.2 | 82.76 |
| Nomic v1 | Smooth K2 Loss | 2.5 | 0.2 | 82.79 |
| Nomic v1 | Smooth K2 Loss | 2 | 0.2 | 82.82 |
| Nomic v1 | Smooth K2 Loss | 3 | 0.2 | 82.83 |
| Jina v2 | Smooth K2 Loss | 3 | 0.15 | 83.51 |
| Jina v2 | Smooth K2 Loss | 3 | 0.2 | 83.54 |
| Jina v2 | Smooth K2 Loss | 3.5 | 0.2 | 83.55 |
| Jina v2 | Smooth K2 Loss | 4 | 0.2 | 83.55 |
The experimental results from Table 5 reveal minor fluctuations in model performance across diverse hyperparameter configurations, which affirms the resilience and robustness of our proposed methodology. This stability highlights the inherent adaptability of our regression framework as well as loss functions, suggesting their applicability to a wide range of modeling scenarios without necessitating extensive hyperparameter optimization.
@@ 4.5 Ablation Studies
In section 4.1, we initially demonstrated the effectiveness of our regression framework by comparing the performance of models utilizing both classification-based and regression-based strategies for multi-category STS tasks. Then, we elucidated the significance of zero-gradient buffer zones by evaluating the performance of models when selecting either Translated ReLU or L1 Loss, and Smooth K2 Loss or MSE Loss as the loss function. These comparisons directly align with the three core innovations of this paper and fulfill the role of ablation experiments.
Here, we extend our ablation study by evaluating our network architecture, as depicted in Figure 1. Specifically, we seek to determine the necessity of concatenating , , and their element-wise difference in the final linear layer of the model. To this end, we employ both BERT and RoBERTa under the same experimental conditions outlined in section 4.1, with the results presented in Table 6. The findings indicate that the concatenation method is the most effective for both PLMs, thus further validating the rationale behind our proposed scheme.
@@ 5 Conclusion
In this paper, we propose an innovative regression framework and develop two simple yet efficacious loss functions: Translated ReLU and Smooth K2 Loss, to address multi-class STS tasks. Compared to traditional classification approaches, our regression modeling strategy effectively captures the progressive relationships between categories, thereby achieving superior performance while reducing the the parameter count in the model’s output layer.
Further empirical evidence demonstrates that our method can also be combined with leading contrastive learning models, leveraging fine-grained annotated data to further enhance their performance. Moreover, this approach proves to be more advantageous than continued fine-tuning through contrastive learning, both in terms of performance gains and computational efficiency.
To support further research, we have made our code and model checkpoints publicly available.
@@
Due to the lack of baselines and computational resource constraints, the experiments in this paper primarily focus on encoder-only discriminative models, rather than recently advanced generative pre-trained models (e.g. LLaMA (Touvron et al., 2023), Mistral (Jiang et al., 2023)). However, it is important to emphasize that, compared to mainstream generative PLMs, the models we selected—BERT, RoBERTa, Jina Embeddings v2, and Nomic Embed v1—have significantly fewer parameters. This results in higher inference efficiency, which is particularly advantageous in large-scale information retrieval and text clustering scenarios.
Appendix A Data Filtering Method
As mentioned in section 4.2, before applying the STS-B and SICK-R training sets for model updates, we implemented strict data filtering to ensure that no sentence pairs present in the test sets would appear in the fine-tuning corpus.
To elaborate on this process, we first take the SICK-R dataset as an example to illustrate the standard format of STS datasets. As shown in Table 7, each sample consists of two text strings, "sentence 1" and "sentence 2," along with a floating-point number "score" that indicates the semantic similarity between them. We denote these as "s1," "s2," and "r," respectively.
Then, for any sentence pair within the STS-B or SICK-R training set, if a sample exists in the test sets of STS12-16, STS-B, or SICK-R such that and , or and , we treat them as duplicates and remove the corresponding sentence pair from the training data. It should be noted that the entire process is conducted without any modifications to the test sets.
This filtering mechanism is stringent, as we do not take into account whether and are equal. In other words, as long as a sentence pair appears in both the training and test sets, it is removed from the training corpus, regardless of whether the similarity scores are identical. Under this protocol, even within the SICK-R dataset itself, there are instances where samples from the training and test sets overlap. Examples in Table 7 illustrate such cases. The goal of this approach is to maximize the model’s generalization ability.
| sentence 1 | sentence 2 | score |
| Sentence pairs in the SICK-R training set | ||
| A man in a blue jumpsuit is courageously performing a wheelie on a motorcycle | The man is doing a wheelie with a motorcycle on ground which is mostly barren | 4.1 |
| The tan dog is watching a brown dog that is swimming in a pond | A pet dog is standing on the bank and is looking at another brown dog in the pond | 4.3 |
| Sentence pairs in the SICK-R test set | ||
| The man is doing a wheelie with a motorcycle on ground which is mostly barren | A man in a blue jumpsuit is courageously performing a wheelie on a motorcycle | 3.7 |
| A pet dog is standing on the bank and is looking at another brown dog in the pond | The tan dog is watching a brown dog that is swimming in a pond | 3.6 |